OpenStudy (anonymous):

If A:B=2:3 and B:C=4:5 and 5A:3C is --------- ?

OpenStudy (ganpat):

so A/ B * B/C = 2/3 * 4 /5 A/C = (2 * 4)/ (3 * 5) 5A/3C = (5 * 2 * 4) / (3 * 3 * 5 ) = 8/9..